Calculate the angular momentum (in kg · m2/s) of Mercury in its orbit around the Sun. (The mass of Mercury is 3.300 ✕ 1023 kg, the orbital radius is 5.790 ✕ 107 km and the orbital period is 0.241 y.)

9.213e32 is wrong
Answer #1

:امك m: 3.34102kg ; Y = 5.79x.107 km I= 0.247 -0.241365x24X60X60 =76001765 I = 76001765._—.-- Angular Velocity@= 27/1 = 27/76